### v2

# Negotiative Alignment: Embracing Disagreement to Achieve Fairer Outcomes – Insights from Urban Studies

Abstract
Cities are not monolithic; they are arenas of negotiation among groups that hold varying needs, values, and experiences. Conventional methods of urban assessment—from standardized surveys to AI-driven evaluations—frequently rely on a single consensus metric (e.g., an average measure of inclusivity or safety). Although such aggregations simplify design decisions, they risk obscuring the distinct perspectives of marginalized populations. In this paper, we present findings from a community-centered study in Montreal involving 35 residents with diverse demographic and social identities, particularly wheelchair users, seniors, and LGBTQIA2+ individuals. Using rating and ranking tasks on 20 urban sites, we observe that disagreements are systematic rather than random, reflecting structural inequalities, differing cultural values, and personal experiences of safety and accessibility.
Based on these empirical insights, we propose negotiative alignment, an AI framework that treats disagreement as an essential input to be preserved, analyzed, and addressed. Negotiative alignment builds on pluralistic models by dynamically updating stakeholder preferences through multi-agent negotiation mechanisms, ensuring no single perspective is marginalized. We outline how this framework can be integrated into urban analytics—and other decision-making contexts—to retain minority viewpoints, adapt to changing stakeholder concerns, and enhance fairness and accountability. The study demonstrates that preserving and engaging with disagreement, rather than striving for an artificial consensus, can produce more equitable and responsive outcomes in urban design.
Keywords: Negotiative alignment, urban design, AI alignment, multi-stakeholder preferences, fairness, accountability, participatory, inclusivity
Introduction
Urban environments appear to serve everyone, yet people experience cities through different social, cultural, and political lenses. A wheelchair user in Montreal may perceive certain sidewalks as insurmountable, while an LGBTQIA2+ resident might interpret street murals or local businesses as signaling inclusion or exclusion. However, contemporary urban analysis tends to emphasize aggregated metrics: a single “overall rating” for aesthetic quality, accessibility, or safety (Mehta 2014; Varna and Tiesdell 2010). Although helpful for city-wide planning, these averages may dilute the specific, and often urgent, needs of marginalized groups (Fainstein 2010; Costanza-Chock 2020).
Drawing on a multi-phase study of public space in Montreal, we show that apparent disagreements in urban evaluations are not merely noisy outliers. Rather, they reflect differences in lived experience, structural inequality, and cultural values. Participants in our study included 35 individuals from multiple demographic categories, such as wheelchair users, seniors, LGBTQIA2+ individuals, and immigrants. By combining semi-structured interviews, rating tasks, and forced-ranking experiments, we found that disagreements were pervasive and persistent, particularly concerning dimensions like accessibility, inclusivity, and practicality. As discussed in Section Results, such divergent viewpoints could not be resolved merely by factual clarifications; instead, they pointed to deeper identity-based needs.
Conventional AI or survey-based methods prioritize consensus by aggregating heterogeneous preferences into a single output score. This approach often obscures critical minority viewpoints (Gabriel 2020; Rawls 1999), such as the need for comprehensive curb cuts for wheelchair users or the importance of affirming cultural symbols for LGBTQIA2+ communities (Anttiroiko and de Jong 2020). While recent advances in AI alignment incorporate diverse perspectives through multi-agent modeling, game-theoretic approaches, and fairness-aware optimization (Zhi-Xuan et al. 2024; Sorensen et al. 2024; Jain et al. 2024; Dwork et al. 2012; Ferrara 2023), they rarely include explicit mechanisms for negotiating conflicting preferences in real time. Addressing this gap requires alignment strategies that iteratively adjust stakeholder weights and trade-offs rather than statically aggregating preferences.
To bridge this gap, we introduce negotiative alignment, a dynamic, multi-agent approach that integrates disagreements into iterative AI decision-making processes (Section Negotiative Alignment Framework). Unlike singular consensus methods, negotiative alignment systematically identifies persistent conflicts, assesses whether they can be partly resolved through design modifications, and updates stakeholder-specific weighting factors to reflect each group’s preferences. By embedding formal bargaining mechanisms (Nash 1950; Kalai and Smorodinsky 1975) in the design pipeline, negotiative alignment aims to enhance fairness and inclusivity. Far from viewing divergence as an anomaly, it treats disagreement as a resource for building more equitable outcomes.
Contribution and Research Questions
Our study reveals why persistent disagreement may be essential for fair and accountable urban planning. We frame our contributions around four core research questions:
  • (RQ1) Who disagrees about what, and why? We identify which design criteria, such as accessibility or inclusivity, yield the strongest disagreements and investigate the structural, cultural, and personal factors underlying them.
  • (RQ2) How do different evaluation methods capture disagreement, particularly among minority groups with intersectional identities? While ranking and rating methods are well-established in social choice theory for aggregating preferences (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023; Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023), we seek to understand how these methods reflect the agreement and disagreement dynamics within minority populations possessing intersectional identities. This focus allows us to examine whether traditional evaluation techniques adequately capture the nuanced preferences and potential conflicts that arise from overlapping social categories such as race, gender, and socioeconomic status.
  • (RQ3) How does negotiation occur in group settings, and can it be systematized? We examine whether focus-group discussions reduce or merely mask disagreements and discuss how these deliberations could inspire algorithmic negotiation procedures.
  • (RQ4) How can we incorporate negotiative alignment into AI models? We propose a multi-agent framework that integrates dynamic preference updates, thereby enabling AI systems to preserve minority viewpoints and adapt to evolving stakeholder needs.
These questions guide the paper’s structure and our interpretation of data. The remainder of the paper is organized as follows. Section 2 reviews related urban planning and AI alignment literature. Section 3 outlines the methodology for our multi-phase study in Montreal. Section 4 presents the results of both rating and ranking tasks, illustrating where and why disagreements arise. Section 5 situates these empirical findings within a proposed negotiative alignment framework, articulating its potential for broader application. We conclude by reflecting on limitations and directions for future research.
Related Literature
Urban assessment frameworks commonly aggregate feedback into a single metric for decision-making (Mehta 2014; Varna and Tiesdell 2010; Naik et al. 2014; Qian et al. 2023; Yurrita et al. 2021). Although efficient for policy or administrative planning, these techniques do not always capture the nuanced needs of marginalized communities (Fainstein 2010; Sandercock 1998; Gowaikar et al. 2024). Inaccessible sidewalks (Imrie 2012) and queer space erasure (Doan 2011) exemplify how aggregate metrics may obscure issues relevant to minority groups. Our findings align with these critiques, suggesting that demographic-specific experiences can yield systematic disagreements worthy of deeper consideration.
Urban planners have historically relied on frameworks such as the Gehl Institute’s public life tools (Gehl and Svarre 2013) and the “Place Diagram” by Project for Public Spaces (Santos Nouri and Costa 2017), as well as indices like the Public Space Index (PSI) (Mehta 2014) and the Pedestrian Environment Quality Index (PSEQI) (Zamanifard et al. 2019) to evaluate aspects of comfort, access, and sociability. While these approaches incorporate observational data and community engagement, they typically yield composite metrics that may mask diverse and context-specific experiences. Recent scholarship in critical urban studies highlights the limitations of consensus-based measures, particularly their potential to underrepresent the needs of marginalized groups. Our work extends this critique by proposing a framework that employs participatory methods to formalize disagreement through a negotiative lens, thereby complementing established approaches and advancing discussions on equity in urban design.
AI Alignment and Pluralistic Models
Traditional AI alignment generally involves mapping a model’s objective function to a representative user preference or a set of aggregated preferences (Gabriel 2020; Rahwan et al. 2019). This approach works adequately for many applications, but it often overlooks persistent disagreements and minority voices. Pluralistic alignment methods (Zhi-Xuan et al. 2024; Sorensen et al. 2024; Nayak et al. 2024) have begun to address this limitation by modeling a distribution of perspectives rather than a single consensus. However, these techniques usually stop short of incorporating active negotiation or formal bargaining procedures to handle direct conflicts between stakeholder groups.
Multi-Agent Reinforcement Learning and Game Theory
Multi-Agent Reinforcement Learning (MARL) extends single-agent paradigms by assigning separate objectives to each agent, sometimes using equilibrium concepts or Pareto-efficiency to mediate conflicts (Zhang, Yang, and Başar 2021; Yang et al. 2020). Game-theoretic approaches, including Nash bargaining (Nash 1950; Nakamura 2024), Kalai-Smorodinsky fairness (Kalai and Smorodinsky 1975; Chen and Hooker 2023), and cooperative resource allocation (Rosen 1965; Doostmohammadian et al. 2025), also provide structured ways to reconcile divergent interests. While these models facilitate coordination among agents with different goals, they typically assume a static set of fairness constraints and do not dynamically adjust stakeholder influence as new information or disagreements arise.
Deliberative AI and Fairness-Conscious Optimization
Deliberative AI focuses on argumentation and structured dialogue to find acceptable outcomes (Walton, Reed, and Macagno 2009; Rahwan and Simari 2009; Nakamura 2024; Fish et al. 2023). Although such approaches emphasize reasoning about conflicting arguments, they often lack iterative mechanisms to update weights across multiple negotiation rounds. Similarly, fairness-conscious optimization seeks to maintain acceptable performance across demographic groups (Moulin 2004; Dwork et al. 2012; Nakamura 2024), but typically imposes static fairness constraints. Disagreement persists if preferences evolve or if new stakeholder groups become relevant.
Negotiative Alignment
To handle disagreements that neither fade over time nor converge under static fairness criteria, negotiative alignment incorporates dynamic bargaining mechanisms. This extension is particularly relevant in urban planning, where stakeholder needs can shift due to evolving infrastructure, demographics, or cultural contexts. Past research in multi-agent urban planning simulates stakeholders with distinct objectives (Qian et al. 2023; Yurrita et al. 2021; Zhou et al. 2024; Ni, Wang, and Liu 2024), but it rarely integrates formal negotiation procedures to address deep-seated conflicts. Dissent-aware learning methods likewise highlight the importance of preserving minority viewpoints (Gordon et al. 2022; Chen, Weld, and Zhang 2021), yet they do not specify how to arbitrate among directly opposing interests once disagreements are identified.
Our approach complements these lines of research by developing a formal negotiation operator that iteratively updates stakeholder weights and preference distributions. This operator ensures that persistent disagreements remain explicit in both the design process and the resulting AI recommendations, potentially strengthening fairness and accountability in domains beyond urban planning.
Methodology
Study Setting: Montreal, Summer 2023–Autumn 2024
We conducted this study in Montreal, a multilingual city characterized by diverse architectural styles and intersecting social communities. This diversity enabled recruitment from multiple marginalized populations, thereby illustrating the variety of urban experiences and how conflicting preferences necessitate a framework for negotiative alignment (Margier 2013; Sylvestre 2010; Mushkani et al. 2025b).
Figure 1 summarizes the self-reported age distribution and intersects additional demographic characteristics. This representation underscores that a participant’s identity does not wholly determine their perceptions but often shapes experiential factors contributing to disagreements in urban evaluations.
Figure 1: Representation of diversity across age groups. The distribution of self-reported demographic characteristics varies by age, highlighting intersectional perspectives and reminding us that no single identity alone defines an individual’s experience (Mushkani et al. 2025b).
Participant Recruitment and Profiles
We contacted over 100 local organizations and established partnerships with more than 30 institutions serving wheelchair users, seniors, LGBTQIA2+ individuals, and recent immigrants (IRCGM 2018; Mushkani et al. 2025b). Although including a majority-resident control group might have offered broader contrasts, our focus was on systematically marginalized participants, whose perspectives are often obscured by aggregate metrics. As a result of these partnerships, 35 individuals participated in the study, with several contributing to multiple phases of data collection:
  • Total participants: 35
  • Overlap across tasks:
    • 28 in-depth semi-structured interviews
    • 12 rating participants (individual and group tasks)
    • 17 ranking participants (individual tasks)
Multi-Phase Data Collection
Phase I: Semi-Structured Interviews
We conducted 28 semi-structured interviews focusing on participants’ conceptualizations of quality in public spaces. Each interview lasted between 40 and 90 minutes and was audio recorded. Questions included how participants use public spaces, how they define and talk about them, which ones are their favorites, and what kinds of experiences they recall when discussing an ideal public space. Additionally, we explored what qualities participants believe an ideal public space should possess. These initial questions were followed by more contextual and in-depth inquiries (Creswell and Creswell 2022).
Phase II: Focus Groups and Rating
Three focus-group sessions, each consisting of 3–5 participants, were organized based on participants’ availability and schedules. During these sessions, individuals first independently rated 20 selected street images on a 4-point scale (1,=,poor; 4,=,excellent) according to four core criteria: inclusivity, accessibility, aesthetics, and practicality. Our selection of 20 street views adhered to a structured socio-spatial matrix (see Appendix Figures 11–13), ensuring representational diversity in land-use, historical contexts, and green-space distribution. Following the individual ratings, participants engaged in group discussions to negotiate their ratings, exploring whether disagreements were due to objective observations (e.g., unrecognized curb ramps) or deeper value differences (e.g., differing notions of safety) (Margier 2013).
Since participants self-selected their groups, we aimed to minimize interference in this process. However, our objective was to ensure diversity within each group and to avoid scenarios where a single identity group was predominant.
Phase III: Ranking Experiment
Seventeen participants were tasked with individually ranking seven images from best to worst based on each of the 12 previously identified criteria. This process resulted in twelve sets of rankings per participant, each corresponding to a different criterion. For each set, Kendall’s Tau was computed to measure the level of agreement among participants regarding the positioning of the images. The ranking exercise compelled participants to make explicit trade-offs, thereby revealing underlying priorities and values influencing their evaluations of urban spaces (Sylvestre 2010). Due to the nature of this task, we reduced the number of images from 20 to 7, allowing us to increase the number of criteria and compare the performance of various concepts more effectively.
Our partnership with community organizations remains ongoing, emphasizing a co-production philosophy through workshops and knowledge exchange. Participants were compensated for their time and contributions, while we also fostered a reciprocal relationship in which participants learned about AI and expert knowledge on cities from the researchers, while we gained insights into public space usage and perceptions from the citizens (IRCGM 2018; Mushkani et al. 2025a). Furthermore, all study procedures were reviewed and approved by the Research Ethics Board, ensuring that ethical standards were maintained throughout the research process.
Analysis Methods
Qualitative Analysis.
We thematically coded interview and focus-group transcripts, emphasizing statements that elucidated the underlying why of disagreements (e.g., cultural symbolism, mobility constraints) (Creswell and Creswell 2022). From the interviews, we identified twelve criteria that were used for the ranking task: accessibility, invitingness, comfort, regenerativeness, beauty, practicality, maintenance, inclusivity, dynamism, representativeness, oppression, and security.
For the rating task, participants prioritized four key criteria: accessibility, inclusivity, aesthetics, and practicality. These criteria were consistently emphasized in discussions as the most relevant for evaluating public spaces, guiding both individual and group assessments.
Quantitative Analysis.
  • Pearson’s Correlation (Puth, Neuhäuser, and Ruxton 2015): We calculated agreement levels between individual and group ratings, as well as between different demographic groups. This analysis assessed the extent to which group discussions influenced rating shifts, identifying cases where collective deliberation significantly altered individual assessments.
  • Kendall’s Tau (Abdi 2006): We measured ordinal ranking similarities across demographic groups for each criterion to identify patterns of agreement and areas of persistent disagreement. This helped pinpoint criteria where differences in weighting were irreconcilable, indicating fundamental divergences in preferences.
These analyses collectively uncovered both reconcilable and persistent disagreements across participant groups, supporting our proposal for a negotiative alignment method capable of preserving minority perspectives rather than averaging them out.
Results
Overview of Street-Level Findings
Figure 2 summarizes the aggregated rating data for 20 Montreal streets, assessed across multiple dimensions. Persistent disagreements often clustered along demographic lines, reinforcing the necessity of preserving disagreements in AI frameworks.
Figure 2: Matrix of compiled ratings for 20 Montreal streets across multiple dimensions (e.g., inclusivity, accessibility, aesthetics, practicality). Ratings are disaggregated by core demographic groups (each group consists of 2 persons), illustrating distinct disagreements.
Illustrative examples include:
  • Rue Rene Goupil (i08). It received a low aggregate score (1.56), driven largely by low inclusivity (1.0) ratings across multiple groups.
  • Place Armes (i16). Wheelchair users scored it at 1.0 for accessibility due to frequent curb obstructions, whereas others rated accessibility higher.
  • Laurier Ave (i14). Most participants viewed it favorably for its aesthetics and inclusivity, but those using wheelchairs criticized its lack of accessibility.
  • Rue du Boree (i06). A wheelchair user gave it an accessibility score of 4.0, but LGBTQIA2+ participants rated its aesthetics at 1.0, citing concerns about symbolic inclusivity and perceived safety.
Such divergences highlight how a single mean metric (e.g., 2.5 out of 4) can mask critical user experiences.
Who Disagrees and Why
Wheelchair User Discrepancies
Wheelchair users diverged from other participants on accessibility and practicality. For example, Rue Commune (i04) was rated positively for aesthetics (3.0) and accessibility (4.0) by some LGBTQIA2+ participants but received a score of 1.0 for both accessibility and aesthetics from wheelchair users, highlighting tangible barriers that were overlooked by those not reliant on mobility devices. For more details, please see the appendix.
LGBTQIA2+ Aesthetic Divergence
LGBTQIA2+ participants often favored symbolic inclusivity (e.g., flags, safe-space signage) and lively street atmospheres. Seniors tended to rate these same environments lower, citing concerns about noise and a sense of overcrowding. Such disagreements underscore that aesthetic judgments often embed cultural and safety considerations, rather than mere personal taste.
Function vs. Appeal
Quantitative correlations showed near-zero or negative relationships between practicality and aesthetics. Cobblestone streets exemplify this trade-off, with high visual appeal but limited wheelchair mobility. The forced-ranking tasks magnified these tensions, requiring participants to prioritize certain dimensions over others.
Post-occupancy vs. Pre-occupancy
Several participants noted that judgments of inclusivity and aesthetics are informed by post-occupancy reporting, including lived experiences and subcultural knowledge. For example, some observed that public art might evoke wariness among seniors and potentially contribute to gentrification, contrasting with its role in beautification as perceived in Montreal’s Mile End neighborhood. Similarly, wheelchair users familiar with specific ramps or tactile paving rated these design elements more favorably due to their enhancement of mobility. This evidence highlights the role of familiarity in shaping disagreements over inclusivity and aesthetics, reinforcing our broader claim that such divergences are systematically rooted in identity and social context rather than individual preference.
Figure 3: Violin Plot of Inclusivity Ratings: Distribution of inclusivity ratings for 20 streets across demographic groups, highlighting differences in perceived inclusivity and the influence of features such as seating and sidewalk conditions.
Negotiation in Group Settings
Group-based discussions partially resolved conflicts stemming from objective misunderstandings (e.g., the presence of ramps) or from evaluations adopting a more needs-based perspective. Engaging in discussions and listening to personal experiences allowed participants to clarify certain discrepancies. However, value-based disagreements rooted in identity or cultural experiences persisted. Negotiation thus helped distinguish whether disputes arose from incomplete information or from fundamentally divergent perspectives on space usage and symbolism.
Rating vs. Ranking: Contrasting Insights
  • Rating. The 4-point scale allowed for partial convergence on less contested dimensions but often masked deeper disagreements. This finding aligns with social choice theory, where rating methods capture the intensity of preferences but may overlook nuanced conflicts (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023).
  • Ranking. Forced-order exercises exposed sharp conflicts when participants were required to assign a single linear priority. For instance, disagreements between accessibility and aesthetics became more pronounced. This observation corroborates previous studies that indicate ranking methods can reveal underlying conflicts more effectively than rating methods (Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023).
Although negotiation could clarify some misunderstandings, the data reveal that certain trade-offs persist. This pattern supports an approach that retains disagreements in computational models instead of converging on an artificial consensus, consistent with recommendations in the social choice literature (Fritz, Moretti, and Staudacher 2023).
Summary of Key Observations
  1. 1.
    Disagreement is Structurally Patterned. Divergences in urban evaluations are not random but systematically reflect structural inequities, particularly for marginalized groups (e.g., mobility, symbolic safety) (Zhu, Zhang, and Biljecki 2025; Adams and Laursen 2007; Cruz, Roskamm, and Charalambous 2018; Mela 2014; Leroy et al. 2024; Jackson, Williams, and VanderWeele 2016; Mushkani and Ono 2021). These findings align with intersectional urban studies, demonstrating that public space assessments are shaped by overlapping exclusions based on race, gender, disability, and socioeconomic status. Our study extends this literature by formalizing these divergences within a computational framework for AI alignment.
  2. 2.
    Ratings and Rankings Capture Distinct Information. Ratings offer finer granularity, allowing for the representation of intersectional positionalities, whereas rankings impose explicit trade-offs, often simplifying multi-dimensional concerns (Xu et al. 2024; Rofin et al. 2023). While prior research has conceptually distinguished these methods, our findings provide empirical validation of their respective strengths and limitations, highlighting the need for hybrid approaches in preference alignment.
  3. 3.
    Negotiation Partially Mitigates Discrepancies. While no single urban site received unanimous evaluation, negotiation facilitated a more nuanced and balanced assessment. As shown in the violin plot (3), negotiative evaluations of urban inclusivity exhibit calibration effects at both individual and group levels. Objective clarifications help refine understanding, yet fundamental value-based disagreements persist across demographic lines, reinforcing the notion that negotiation alone cannot fully resolve underlying value conflicts (Suzuki and Horita 2024). This dynamic reflects the inherent nature of cities as arenas of continuous negotiation among diverse individuals and entities (Lefebvre 1974; Jacobs 1961; Harvey 1989).
  4. 4.
    Averages Overlook Marginalization. A mean rating can obscure experiences of groups who find a space unusable or exclusionary, reinforcing concerns about aggregative methods masking minority viewpoints (Fritz, Moretti, and Staudacher 2023).
Our study not only corroborates existing findings in social choice theory regarding the strengths and limitations of rating and ranking methods but also extends these insights by highlighting how intersectional identities influence agreement and disagreement dynamics. This underscores the necessity of incorporating minority perspectives to achieve a more comprehensive and equitable understanding of stakeholder preferences in urban planning.
Further Analyses
Figures 4 through 8 present correlation metrics (Pearson’s and Kendall’s Tau) across multiple criteria (e.g., accessibility, inclusivity, aesthetics, practicality, security). These visualizations demonstrate how disagreements vary among demographic groups and highlight the need for an approach that preserves diversity in preferences:
Figure 4: Rating using Pearson correlation coefficients between demographic groups for practicality, inclusivity, aesthetics, and accessibility. Higher disagreement is observed in accessibility and practicality scores, particularly among participants with intersectional identities. Notably, the individual identifying as female, Black, senior, and disabled, as well as the individual identifying as female, minority, adult, and disabled, exhibit greater divergence from the rest.
Figure 5: Ranking using Kendall Tau coefficients for accessibility, invitingness, comfort, and regenerativeness. Certain criteria, such as regenerativeness, show higher consensus, whereas accessibility remains more contentious.
Figure 6: Ranking using Kendall Tau for beauty, practicality, maintenance, and inclusivity. Disagreement spikes in practicality and inclusivity, reflecting differences in lived experiences.
Figure 7: Ranking using Kendall Tau for dynamism, representativeness, oppressiveness, and security. These dimensions often polarize participants, reflecting divergent perceptions of safety, vibrancy, and cultural representation.
Figure 8: Group Rating using Pearson correlation coefficients for selected citizen groups (F1G, F2G, F3G) across practicality, inclusivity, aesthetics, and accessibility. The results highlight a more balanced evaluation.
Negotiative Alignment Framework
Our findings corroborate existing social choice theory regarding the strengths and limitations of rating and ranking methods (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023; Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023). Specifically, ratings captured the intensity of preferences, aligning with theories that emphasize their ability to reflect nuanced stakeholder values, while rankings effectively exposed fundamental conflicts, as previously documented. Beyond these confirmations, our study extends the literature by demonstrating how minorities with intersectional identities shape the dynamics of agreement and disagreement. For instance, wheelchair users and LGBTQIA2+ participants exhibited distinct preferences that traditional aggregation methods might overlook (Leroy et al. 2024; Xu et al. 2024). This highlights the necessity of incorporating negotiative alignment into AI models, which treats these persistent disagreements as meaningful inputs rather than statistical noise. By doing so, AI systems can better preserve minority viewpoints and adapt to the complex, evolving needs of diverse urban populations (Zhou et al. 2024; Kirk et al. 2024; Sorensen et al. 2024; Mushkani et al. 2025b).
Multi-Agent Utility and Constraint Formulation
Let $\mathcal{X}$ be a set of possible urban design options, and let $G$ index stakeholder groups (e.g., wheelchair users, LGBTQIA2+ communities). Each group $g$ is characterized by:
$P_{g}:\mathcal{X}\to\mathbb{R},$
which quantifies that group’s preference for each option in $\mathcal{X}$. Traditional models maximize:
$\max_{x\in\mathcal{X}}\sum_{g=1}^{G}\lambda_{g}P_{g}(x),$
subject to $\sum\lambda_{g}=1,\,\lambda_{g}\geq 0$. This averaging tends to underrepresent minority experiences (Rawls 2001).
Dynamic Multi-Agent Negotiation
In negotiative alignment, weights $\lambda_{g}(t)$ are updated iteratively:
$\lambda_{g}(t+1)=f(\{\lambda_{h}(t)\},\{\Delta_{g}(t)\},C),$
where $\Delta_{g}(t)$ captures newly surfaced disagreements at iteration $t$, and $C$ includes resource constraints or policy requirements (Moulin 2004). The system re-optimizes after each iteration, reflecting evolving stakeholder concerns.
Iterative Negotiation Operator
We introduce a negotiation operator $\mathcal{N}$:
$(x^{*}(t+1),\{\lambda_{g}(t+1)\})=\mathcal{N}((x^{*}(t),\{\lambda_{g}(t)\}),\{P_{g}\}),$
which identifies critical disagreements, proposes possible design or weighting adjustments, and employs a rule such as Nash or Kalai-Smorodinsky bargaining (Nash 1950; Kalai and Smorodinsky 1975) to guide solutions.
Identity Preservation and Fairness Metrics
To avoid erasure of minority preferences, we define an identity-preserving preference distribution:
$\pi_{g}(t+1)=\beta\,\pi_{g}(t)+(1-\beta)\,\hat{\pi}_{g},$
where $\hat{\pi}_{g}$ is the updated preference distribution, and $\beta$ controls how quickly preferences adapt (Zinkevich 2003).
We also introduce an Identity Preservation Index (IPI) using Jensen-Shannon divergence:
$\text{IPI}(g,t)=1-D_{JS}\bigl{(}\pi_{g}^{0}\,\|\,\pi_{g}(t)\bigr{)},$
which measures how closely current preferences align with the original distribution $\pi_{g}^{0}$. A high IPI indicates that a group’s foundational needs remain visible in the final decision.
Benchmarks for Evaluating Negotiative Alignment
We propose the following metrics:
Disagreement Coverage Ratio (DCR).
$\text{DCR}=\frac{\sum_{g=1}^{G}\mathbb{I}\left(\exists\,x:|P_{g}(x)-P_{g^{\prime}}(x)|\geq\delta\right)}{G},$
where $\delta$ is a threshold for disagreement. A higher DCR indicates that the model is preserving significant group-specific disagreements rather than collapsing them into an average.
Negotiation Progress Metric (NPM).
$\text{NPM}(t)=\frac{1}{G}\sum_{g=1}^{G}\|\lambda_{g}(t+1)-\lambda_{g}(t)\|_{1}.$
This metric tracks how much the weighting of different groups shifts between iterations. Non-zero changes indicate that the system is actively adjusting to stakeholder feedback.
Prototype Demonstration
To illustrate negotiative alignment, we implemented a small-scale prototype on subset Vivian Avenue (i09) (inclusivity dimension). Table 1 shows each group’s rating, and Table 3 presents the iterative consensus after two negotiation rounds. This serves as a microcosm of the full analysis across 20 urban scenes (Appendix A7). Our Python code (Appendix A6) updates group weights based on disagreement magnitudes while preserving identity distributions via a Jensen–Shannon-based metric.
LGBTQIA2+ Handi-capped Elderly Female Elderly Male Young Male Young Female
3.0 2.0 1.7 2.0 2.0 1.7
Table 1: Stakeholder ratings for urban scene i09.
The prototype uses an iterative negotiation operator that updates the consensus rating and group weights based on absolute disagreement. Key metrics include the Disagreement Coverage Ratio (DCR), the Negotiation Progress Metric (NPM), and the Identity Preservation Index (IPI). Iteration continues until weight changes fall below 0.01 or after 10 iterations. Table 2 summarizes the initial iterations:
Iteration Weights
0 [0.167, 0.167, 0.167, 0.167, 0.167, 0.167]
1 [0.333, 0.101, 0.182, 0.101, 0.101, 0.182]
2 [0.322, 0.096, 0.196, 0.096, 0.096, 0.196]
Table 2: Weight adjustments across iterations in the negotiation process.
Iteration $x^{*}$ DCR (%) Avg IPI (%)
0 2.067 16.7 90.5
1 2.224 50.0 85.6
2 2.204 50.0 84.0
Table 3: Consensus value $x^{*}$, Disagreement Coverage Ratio (DCR), and Identity Preservation Index (IPI) during negotiation iterations.
During the empirical study, participants’ group rating for i09 was 2.6. The naive average, computed across all participants, was 2.067, while the negotiative alignment score was 2.224, reflecting a moderate shift toward the group rating. Detailed calculations for all 20 urban scenes—evaluated for inclusivity, accessibility, practicality, and aesthetics—are provided in Appendix A7. These findings illustrate how selecting appropriate hyperparameters and iteration settings can affect the resulting negotiated score. The robustness of the final output depends on both the quantity and specificity of agent-related data, as well as on parameter tuning within the model. This iterative weighting process further clarifies the relative emphasis placed on each group’s viewpoint.
This prototype demonstrates the framework’s capacity to refine consensus while preserving critical elements of each group’s initial preferences. However, it relies on a single criterion and participant demographic data, without additional contextual or agent-specific information. Further evaluation is necessary to verify the framework’s applicability under more comprehensive conditions.
Integrating Large Language Models (LLMs) into the Negotiative Alignment Loop
The negotiative alignment framework can be instantiated with LLMs by treating each stakeholder group as an LLM agent, whose preferences are encoded through prompts or fine-tuning. Each agent receives a summary of the current design proposal and a concise representation of other agents’ positions, then generates a counter-proposal or weight adjustment. Because LLMs have finite context windows, the negotiation protocol must specify which aspects of the design and which past exchanges are carried forward. In practice, this can be achieved by maintaining a rolling buffer of key arguments or by distilling earlier rounds into abstracted summaries that preserve points of contention without exceeding the model’s input limits.
Negotiation proceeds either in an iterative manner—where agents exchange multiple rounds of proposals to approximate middle grounds—or as a one-shot exchange—where all relevant information is compressed into a single prompt. An iterative protocol can mitigate context constraints by revealing information gradually, thereby allowing the model to focus on the most salient disagreements at each step. Conversely, a one-shot protocol may be preferable when the number of stakeholders is small and the full set of arguments fits within the LLM’s input capacity. In both cases, the negotiation operator $\mathcal{N}$ is realized as a sequence of prompt templates that apply formal bargaining rules (e.g., Nash or Kalai–Smorodinsky) to the agents’ generated weight adjustments.
Within each round, LLM agents can apply internal voting mechanisms to generate collective proposals. A majority-voting prompt instructs the model to select the most supported design features, while preference-ranking prompts ask the model to order options by stakeholder-weighted utility. These procedures can be embedded directly into the prompt engineering, enabling the LLM to simulate both deliberation and decision aggregation without external algorithms. The resulting proposals and weight updates are then fed back into the negotiation loop, closing the chain of prompts that implements the iterative operator.
The outputs of each LLM-mediated round also yield text-based records that facilitate the computation of fairness metrics. For example, identity preservation can be measured by comparing the distribution of model-reported preferences at each step against the original set of stakeholder prompts, using divergence measures such as Jensen–Shannon. The degree of negotiation, or the extent of weight adjustment between rounds, can be quantified by tracking the absolute differences in the weight allocations reported by the LLM. By leveraging the generative and reasoning capabilities of large language models, the framework supports a transparent, auditable negotiation process in which disagreement is both preserved in the text transcripts and systematically addressed through successive prompt exchanges.
This paper has articulated the need for more effective negotiation methods in AI-driven decision-making and introduced negotiative alignment as a candidate framework. Future work should implement and empirically evaluate LLM-based negotiation protocols and design studies to assess negotiative alignment in practice. Further research is necessary to advance these methods across diverse application contexts.
Implications for Fairness and Accountability
By allowing disagreements to persist, negotiative alignment frameworks can prevent majority views from overshadowing minority preferences. The iterative updates of $\{\lambda_{g}(t)\}$ produce an auditable record of trade-offs. This transparency can strengthen accountability: when marginalized viewpoints are overridden, stakeholders can identify where and why it happened, and what form of bargaining or design constraint contributed to that decision.
Philosophical and Ethical Considerations
Negotiative alignment views persistent value disagreements as legitimate, rather than anomalies to be averaged out. This resonates with pluralistic conceptions of justice and fairness (Sen 2009). However, practical deployment requires careful attention to how negotiation power is distributed among stakeholders. Mechanism design strategies may help allocate bargaining power more equitably, ensuring no single group unilaterally dictates outcomes.
Limitations and Future Directions
  • Computational Complexity. Solving dynamic multi-agent optimization problems iteratively is computationally demanding, particularly in large-scale urban contexts where numerous variables and constraints must be considered.
  • Real-Time Feedback. Integrating continuous updates from streaming user feedback necessitates the use of efficient reinforcement learning or online optimization algorithms to ensure adaptive and responsive decision-making.
  • Integration with Large Models. Adapting negotiation protocols for large language models or multimodal AI systems presents technical challenges in aligning textual and visual data under evolving preference distributions. Ensuring coherent interpretation and alignment across modalities remains a critical area for future research.
  • Initialization and Trust. The initial weight assignments in preference modeling significantly impact final outcomes. Developing robust initialization strategies that explicitly prioritize the inputs of historically marginalized groups may enhance stakeholder trust and fairness in decision-making processes.
  • Applicability of Negotiative Alignment. Negotiative alignment is particularly suited for multi-stakeholder environments where preferences are dynamic and potentially conflicting. By actively preserving minority perspectives, this approach is well-positioned for applications in urban planning, policy-making, and AI governance.
  • Scope and Sample Limitations. The study’s focus on 35 participants in Montreal captures rich, context-specific insights but may not fully generalize to more diverse urban settings. Expanding to larger or more heterogeneous samples could reveal additional complexities.
  • Feasibility and Real-World Implementation. Real-world application of negotiative alignment requires iterative interactions with multi-stakeholder committees, robust weighting schemes to mitigate power imbalances, and substantial computational resources for repeated updates.
  • Validation and Convergence. Future work could include a co-designed user study with community partners to evaluate how well the negotiative alignment framework resonates with participants’ lived experiences.
  • Structural Inequalities and Power-Sensitive Negotiation. Straightforward bargaining mechanisms (e.g., Nash bargaining) assume a level playing field that does not reflect observed structural inequalities. Future investigations could explore power-sensitive negotiation rules to ensure that historically disadvantaged groups retain sufficient influence throughout the decision-making process.
Conclusion
This study demonstrates how structural inequities, cultural symbols, and personal experiences lead to persistent disagreements when evaluating urban environments. Our findings indicate that conventional consensus-based or averaging methods risk obscuring the very perspectives they aim to serve. We propose negotiative alignment as a framework that preserves such disagreements, dynamically updating model parameters in response to stakeholder negotiations. Empirical evidence suggests that bridging objective misunderstandings can alleviate some disagreements, yet many tensions reflect fundamentally different values. By explicitly modeling these differences, negotiative alignment offers a more robust and ethically grounded approach to designing AI systems for urban and other socially critical domains.
Appendix A Supplementary Material
Heatmap of Correlations by Demographic Group
Figure 9: Correlation heatmap illustrating the relationships between ratings for 20 urban sites across four evaluation criteria—practicality, inclusivity, aesthetics, and accessibility—segmented by demographic groups (LGBTQIA2+, individuals with disabilities, elderly females, elderly males, young males, young females, and group-based evaluations). We observe that practicality and accessibility tend to exhibit stronger correlations within most groups, suggesting some alignment in how these dimensions are evaluated. In contrast, correlations involving aesthetics and inclusivity are generally weaker or negative, possibly reflecting divergent perceptions or trade-offs between these criteria. Group-based evaluations show greater consistency with individual ratings for inclusivity and practicality compared to aesthetics, which appears more variable. Notable differences between demographic groups, such as weaker correlations between accessibility and inclusivity for young males and individuals with disabilities, may indicate varying priorities or lived experiences. These findings suggest the need for AI models to account for diverse perspectives when interpreting such evaluation data.
Criteria Rating Table
Figure 10: Criteria Rating Table: Summary of the four evaluation criteria and their corresponding rating levels. A score of 1 represents the lowest level of performance, while a score of 4 represents the highest. Participants used a sticker-based rating system with color coding to visually represent their assessments. This method allowed for straightforward comparison of ratings across the selected criteria.
Representative Street Views
Figure 11: Representative Street Views: A selection of street-level images from 20 Montreal neighborhoods, capturing variations in land use, pedestrian infrastructure, greenery, and streetscape design. These images provide visual context for the urban spaces included in the study and help to illustrate the diversity of the evaluated sites.
Spatial Distribution of Study Sites
Figure 12: Spatial Distribution of Study Sites: Map of the geographic locations of the 20 selected urban sites across Montreal. We selected these sites to represent a broad spectrum of socio-economic contexts, land uses, urban densities, and historical characteristics. The spatial distribution highlights the diversity of the sites included in our analysis. The base map is sourced from OpenStreetMap.
Street Diversity Classification Guide
Figure 13: Street Diversity Classification Guide: Matrix categorizing the selected streets based on socio-spatial attributes, including land use type, history, greenery, and aesthetic features. Our selection process aimed to ensure diversity by incorporating at least one representative street for each category. For example, streets were chosen to include residential, commercial, and mixed-use areas, as well as variations in history and vegetation levels. This approach provided a structured yet flexible framework for selecting sites that reflect the heterogeneity of Montreal’s urban environment.
Prototype Demonstration Code
1import numpy as np
2import pandas as pd
3from scipy.spatial.distance import jensenshannon
4
5# --- Helper Functions ---
6
7def one_hot(rating):
8 """
9 Convert a scalar rating to a one-hot vector for ratings 1 to 5.
10 Rating is rounded and clipped between 1 and 5.
11 """
12 rating_bin = int(np.clip(round(rating), 1, 5))
13 vec = np.zeros(5)
14 vec[rating_bin - 1] = 1.0
15 return vec
16
17def negotiation_operator(P, lambdas, gamma):
18 """
19 Negotiation operator that updates consensus and stakeholder weights.
20
21 Parameters:
22 - P: Array of group ratings (P_g)
23 - lambdas: Current weights for each stakeholder group
24 - gamma: Step size for weight update (0 < gamma <= 1)
25
26 Procedure:
27 1. Compute the consensus (x*) as the weighted average of P.
28 2. Compute the disagreement Delta for each group as the absolute difference:
29 Delta[g] = |P[g] - x*|.
30 3. Derive a normalized disagreement vector (if total disagreement is zero, use uniform weights).
31 4. Update weights via linear interpolation:
32 lambdas_new = (1 - gamma) * lambdas + gamma * (Delta_normalized)
33 5. Normalize the new weights.
34
35 Returns:
36 x_star, Delta, lambdas_new, and a copy of the previous lambdas (for NPM calculation).
37 """
38 # Step 1: Compute consensus
39 x_star = np.dot(lambdas, P)
40
41 # Step 2: Compute disagreement per group
42 Delta = np.abs(P - x_star)
43
44 # Step 3: Normalize disagreements
45 if np.sum(Delta) == 0:
46 disagreement_normalized = np.ones_like(Delta) / len(Delta)
47 else:
48 disagreement_normalized = Delta / np.sum(Delta)
49
50 # Save old weights for progress metric
51 lambdas_old = lambdas.copy()
52
53 # Step 4: Update weights (simplified bargaining rule)
54 lambdas_new = (1 - gamma) * lambdas + gamma * disagreement_normalized
55
56 # Step 5: Normalize the updated weights
57 lambdas_new = lambdas_new / np.sum(lambdas_new)
58
59 return x_star, Delta, lambdas_new, lambdas_old
60
61def update_identity_preservation(pi_current, x_star, Delta, max_delta=4.0):
62 """
63 Update each group’s preference distribution to preserve its identity when disagreement is high.
64
65 Instead of using the original (static) preference distribution, we update using the current
66 distribution pi_current. We blend the current distribution with the consensus (represented
67 as a one-hot vector from x_star) using a blending factor beta that increases with disagreement.
68
69 Here we set:
70 beta_g = min(Delta[g] / max_delta, 1)
71 so that:
72 - If Delta is high (close to max_delta), beta is near 1 (preserve current identity).
73 - If Delta is low, beta is near 0 (adapt toward consensus).
74
75 Returns:
76 Updated preference distributions for all groups.
77 """
78 # Get consensus as one-hot vector
79 consensus_vec = one_hot(x_star)
80 pi_new = np.zeros_like(pi_current)
81
82 for g in range(len(pi_current)):
83 # Compute disagreement-dependent blending factor
84 beta = np.clip(Delta[g] / max_delta, 0, 1)
85 # Update: blend current identity with consensus
86 pi_new[g] = beta * pi_current[g] + (1 - beta) * consensus_vec
87 # Normalize the updated distribution
88 pi_new[g] = pi_new[g] / np.sum(pi_new[g])
89
90 return pi_new
91
92def compute_IPI(pi_orig, pi_current):
93 """
94 Compute the Identity Preservation Index (IPI) for each group.
95
96 IPI(g) = 1 - (Jensen-Shannon divergence between the original distribution (pi_orig)
97 and the current distribution (pi_current))^2.
98
99 A higher IPI indicates that the group’s updated preference distribution remains closer
100 to its original identity.
101 """
102 jsd = jensenshannon(pi_orig, pi_current, base=2.0)**2
103 return 1 - jsd
104
105# --- Parameters and Initialization ---
106
107# Stakeholder ratings for 6 groups:
108# [LGBTQIA2+, Handicapped, Elderly Female, Elderly Male, Young Male, Young Female]
109P_orig = np.array([3.0, 2.0, 1.7, 2.0, 2.0, 1.7])
110G = len(P_orig)
111
112# Initial weights: uniform distribution across groups
113lambdas = np.ones(G) / G
114
115# Compute the initial consensus rating (x*)
116x_star = np.dot(lambdas, P_orig)
117
118# Create one-hot vectors for the initial preference distributions (pi^0)
119pi_initial = np.array([one_hot(p) for p in P_orig])
120# Save the original (fixed) preferences for IPI calculation
121pi_orig = pi_initial.copy()
122# Start with the current distribution equal to the initial one
123pi_current = pi_initial.copy()
124
125# Negotiation and update parameters
126gamma = 0.5 # Step-size for weight update (negotiation operator)
127epsilon = 0.01 # Convergence threshold for weight changes
128max_iterations = 10
129
130# Fairness metric parameter: threshold for disagreement (for DCR calculation)
131delta_threshold = 0.45
132
133# For storing simulation results
134results = []
135
136# --- Iterative Negotiation Simulation ---
137
138for t in range(max_iterations + 1):
139 # 1. Update consensus and stakeholder weights using the negotiation operator
140 x_star, Delta, lambdas_new, lambdas_old = negotiation_operator(P_orig, lambdas, gamma)
141
142 # 2. Fairness Metrics:
143 # Disagreement Coverage Ratio (DCR): percentage of groups with disagreement >= delta_threshold
144 DCR = (np.sum(Delta >= delta_threshold) / G) * 100
145
146 # Negotiation Progress Metric (NPM): average absolute change in weights between iterations
147 if t == 0:
148 NPM = 0.0
149 else:
150 NPM = np.mean(np.abs(lambdas_new - lambdas_old))
151
152 # 3. Update identity preservation distributions
153 pi_updated = update_identity_preservation(pi_current, x_star, Delta, max_delta=4.0)
154
155 # 4. Compute Identity Preservation Index (IPI) for each group
156 IPI_vals = np.array([compute_IPI(pi_orig[g], pi_updated[g]) for g in range(G)])
157 Avg_IPI = np.mean(IPI_vals) * 100 # Expressed as a percentage
158
159 # Record current iteration metrics
160 results.append({
161 "Iteration": t,
162 "Consensus (x*)": round(x_star, 3),
163 "Weights": [round(w, 3) for w in lambdas],
164 "DCR (%)": round(DCR, 1),
165 "NPM": round(NPM, 3),
166 "Avg IPI (%)": round(Avg_IPI, 1)
167 })
168
169 # Prepare for next iteration: update weights and current distributions
170 lambdas = lambdas_new.copy()
171 pi_current = pi_updated.copy()
172
173 # Check convergence of weight updates (if change is below epsilon, break early)
174 if t > 0 and np.max(np.abs(lambdas - lambdas_old)) < epsilon:
175 break
176
177# Display simulation results
178df_results = pd.DataFrame(results)
179print(df_results.to_string(index=False))
Listing 1: Negotiative Alignment Code for Iterative Stakeholder Weight Updates and Identity Preservation
Table 4: A.7 Simulation Results for Urban Scenes: The final consensus reflects outcomes from applying the framework; empirical consensus is the negotiated agreement among participants during the empirical study. FC: Final Consensus. EC: Empirical Consensus.
PointID Criterion FC Final Weights DCR (%) NPM Avg IPI (%) Group Ratings Naive Average EC
i10 Practicality 3.072 [0.356, 0.065, 0.305, 0.065, 0.065, 0.145] 0.0 0.004 83.4 [3.5, 3.0, 2.7, 3, 3, 2.9] 3.017 2.8
i10 Inclusivity 2.439 [0.042, 0.239, 0.147, 0.239, 0.303, 0.030] 16.7 0.006 66.7 [2.5, 2.0, 2.7, 2, 3, 2.4] 2.433 2.5
i10 Aesthetics 1.734 [0.124, 0.335, 0.255, 0.124, 0.124, 0.039] 33.3 0.003 83.5 [2.0, 1.0, 2.3, 2, 2, 1.8] 1.850 1.5
i10 Accessibility 2.119 [0.121, 0.274, 0.182, 0.350, 0.043, 0.030] 50.0 0.003 50.2 [2.5, 3.0, 2.7, 1, 2, 2.2] 2.233 2.5
i09 Practicality 1.636 [0.192, 0.192, 0.042, 0.192, 0.338, 0.042] 16.7 0.005 83.4 [2.0, 2.0, 1.7, 2, 1, 1.7] 1.733 2.0
i09 Inclusivity 2.240 [0.320, 0.105, 0.220, 0.105, 0.105, 0.144] 33.3 0.004 83.9 [3.0, 2.0, 1.7, 2, 2, 1.9] 2.100 2.6
i09 Aesthetics 2.296 [0.287, 0.173, 0.078, 0.307, 0.078, 0.078] 50.0 0.005 51.4 [3.5, 3.0, 2.0, 1, 2, 2.0] 2.250 2.0
i09 Accessibility 2.279 [0.179, 0.242, 0.142, 0.180, 0.180, 0.078] 100.0 0.005 26.0 [3.5, 4.0, 1.3, 1, 1, 1.8] 2.100 2.0
i06 Practicality 3.266 [0.124, 0.335, 0.255, 0.124, 0.124, 0.039] 33.3 0.003 83.5 [3.0, 4.0, 2.7, 3, 3, 3.2] 3.150 3.0
i06 Inclusivity 2.275 [0.149, 0.176, 0.270, 0.176, 0.176, 0.054] 0.0 0.003 83.3 [2.5, 2.0, 2.7, 2, 2, 2.2] 2.233 2.0
i06 Aesthetics 2.026 [0.161, 0.342, 0.112, 0.325, 0.014, 0.046] 50.0 0.004 66.9 [2.5, 1.0, 1.7, 3, 2, 1.9] 2.017 2.0
i06 Accessibility 2.706 [0.177, 0.326, 0.105, 0.177, 0.177, 0.037] 66.7 0.004 18.8 [2.0, 4.0, 2.3, 2, 2, 2.6] 2.483 1.0
i18 Practicality 2.891 [0.106, 0.323, 0.139, 0.187, 0.187, 0.058] 83.3 0.002 20.1 [3.5, 1.0, 3.7, 4, 4, 3.2] 3.233 2.2
i18 Inclusivity 2.633 [0.166, 0.310, 0.129, 0.124, 0.259, 0.012] 83.3 0.003 34.5 [3.5, 1.0, 3.3, 2, 4, 2.6] 2.733 2.5
i18 Aesthetics 2.442 [0.104, 0.312, 0.271, 0.125, 0.125, 0.063] 66.7 0.005 18.9 [2.0, 1.0, 3.7, 3, 3, 2.7] 2.567 1.5
i18 Accessibility 2.428 [0.233, 0.309, 0.191, 0.101, 0.128, 0.038] 66.7 0.006 35.7 [3.5, 1.0, 3.3, 2, 3, 2.3] 2.517 2.0
i01 Practicality 2.439 [0.042, 0.239, 0.147, 0.303, 0.239, 0.030] 16.7 0.006 66.7 [2.5, 2.0, 2.7, 3, 2, 2.4] 2.433 2.2
i01 Inclusivity 2.409 [0.063, 0.232, 0.068, 0.337, 0.232, 0.068] 16.7 0.006 83.4 [2.5, 2.0, 2.3, 3, 2, 2.3] 2.350 2.3
i01 Aesthetics 2.980 [0.222, 0.222, 0.067, 0.211, 0.211, 0.067] 66.7 0.004 34.6 [4.0, 4.0, 2.7, 2, 2, 2.7] 2.900 3.1
i01 Accessibility 2.233 [0.162, 0.162, 0.325, 0.162, 0.162, 0.028] 16.7 0.002 83.3 [2.0, 2.0, 2.7, 2, 2, 2.2] 2.150 2.1
i17 Practicality 2.000 [0.003, 0.495, 0.003, 0.495, 0.003, 0.003] 33.3 0.003 66.7 [2.0, 3.0, 2.0, 1, 2, 2.0] 2.000 2.0
i17 Inclusivity 1.636 [0.192, 0.192, 0.042, 0.338, 0.192, 0.042] 16.7 0.005 83.4 [2.0, 2.0, 1.7, 1, 2, 1.7] 1.733 2.0
i17 Aesthetics 1.639 [0.299, 0.224, 0.032, 0.224, 0.132, 0.090] 50.0 0.006 50.1 [2.5, 1.0, 1.7, 1, 2, 1.4] 1.600 2.0
i17 Accessibility 2.393 [0.186, 0.286, 0.148, 0.186, 0.186, 0.009] 16.7 0.002 66.7 [2.0, 3.0, 2.7, 2, 2, 2.4] 2.350 2.0
i19 Practicality 2.192 [0.095, 0.352, 0.040, 0.237, 0.237, 0.040] 50.0 0.006 50.9 [2.5, 1.0, 2.3, 3, 3, 2.3] 2.350 3.0
i19 Inclusivity 1.561 [0.042, 0.303, 0.147, 0.239, 0.239, 0.030] 16.7 0.006 66.7 [1.5, 1.0, 1.3, 2, 2, 1.6] 1.567 1.0
i19 Aesthetics 3.266 [0.124, 0.124, 0.255, 0.335, 0.124, 0.039] 33.3 0.003 83.5 [3.0, 3.0, 2.7, 4, 3, 3.2] 3.150 3.0
i19 Accessibility 1.500 [0.005, 0.247, 0.247, 0.247, 0.247, 0.005] 66.7 0.003 40.6 [1.5, 1.0, 1.0, 2, 2, 1.5] 1.500 1.0
i16 Practicality 2.627 [0.084, 0.349, 0.146, 0.084, 0.291, 0.045] 50.0 0.004 69.1 [3.0, 1.0, 3.3, 3, 4, 2.8] 2.850 2.6
i16 Inclusivity 2.350 [0.036, 0.241, 0.172, 0.241, 0.292, 0.019] 66.7 0.006 36.8 [2.5, 1.0, 3.3, 1, 4, 2.3] 2.350 2.3
i16 Aesthetics 2.408 [0.270, 0.350, 0.075, 0.148, 0.148, 0.009] 66.7 0.003 17.3 [3.5, 1.0, 2.7, 3, 3, 2.4] 2.600 1.6
i16 Accessibility 2.259 [0.135, 0.227, 0.259, 0.227, 0.135, 0.016] 83.3 0.002 17.7 [3.0, 1.0, 3.7, 1, 3, 2.2] 2.317 2.1
i11 Practicality 2.240 [0.320, 0.105, 0.220, 0.105, 0.105, 0.144] 33.3 0.004 83.9 [3.0, 2.0, 1.7, 2, 2, 1.9] 2.100 2.2
i11 Inclusivity 1.734 [0.076, 0.182, 0.182, 0.182, 0.313, 0.064] 66.7 0.005 34.3 [2.0, 1.0, 1.0, 1, 3, 1.5] 1.583 1.5
i11 Aesthetics 1.483 [0.216, 0.198, 0.086, 0.198, 0.216, 0.086] 66.7 0.006 67.1 [2.0, 1.0, 1.3, 1, 2, 1.3] 1.433 1.0
i11 Accessibility 2.432 [0.106, 0.376, 0.036, 0.341, 0.106, 0.036] 33.3 0.003 67.7 [2.0, 4.0, 2.3, 1, 2, 2.3] 2.267 2.5
i08 Practicality 1.959 [0.160, 0.302, 0.193, 0.278, 0.017, 0.050] 66.7 0.003 50.2 [2.5, 3.0, 1.3, 1, 2, 1.8] 1.933 1.9
i08 Inclusivity 1.370 [0.076, 0.319, 0.188, 0.188, 0.188, 0.042] 16.7 0.005 66.7 [1.5, 2.0, 1.0, 1, 1, 1.3] 1.300 1.3
i08 Aesthetics 1.000 [0.167, 0.167, 0.167, 0.167, 0.167, 0.167] 0.0 0.000 100.0 [1.0, 1.0, 1.0, 1, 1, 1.0] 1.000 1.3
i08 Accessibility 1.439 [0.042, 0.303, 0.147, 0.239, 0.239, 0.030] 16.7 0.006 50.0 [1.5, 2.0, 1.7, 1, 1, 1.4] 1.433 1.4
i07 Practicality 2.186 [0.288, 0.172, 0.112, 0.172, 0.172, 0.084] 0.0 0.003 100.0 [2.5, 2.0, 2.3, 2, 2, 2.1] 2.150 2.0
i07 Inclusivity 1.814 [0.288, 0.172, 0.112, 0.172, 0.172, 0.084] 0.0 0.003 100.0 [1.5, 2.0, 1.7, 2, 2, 1.9] 1.850 1.0
i07 Aesthetics 2.158 [0.251, 0.251, 0.052, 0.344, 0.052, 0.052] 50.0 0.004 50.3 [3.0, 3.0, 2.0, 1, 2, 2.0] 2.167 2.0
i07 Accessibility 1.617 [0.142, 0.142, 0.250, 0.227, 0.227, 0.012] 50.0 0.002 66.7 [2.0, 2.0, 2.3, 1, 1, 1.6] 1.650 1.0
i02 Practicality 2.507 [0.211, 0.211, 0.089, 0.217, 0.217, 0.054] 66.7 0.004 50.1 [3.0, 2.5, 2.0, 3.0, 3.0, 2.0] 2.517 2.5
i02 Inclusivity 2.186 [0.288, 0.172, 0.112, 0.172, 0.172, 0.084] 0.0 0.003 100.0 [2.5, 2.0, 2.3, 2, 2, 2.1] 2.150 2.5
i02 Aesthetics 2.000 [0.167, 0.167, 0.167, 0.167, 0.167, 0.167] 0.0 0.000 100.0 [2.0, 2.0, 2.0, 2, 2, 2.0] 2.000 2.2
i02 Accessibility 3.233 [0.162, 0.162, 0.325, 0.162, 0.162, 0.028] 16.7 0.002 83.3 [3.0, 3.0, 3.7, 3, 3, 3.2] 3.150 2.5
i14 Practicality 2.667 [0.271, 0.219, 0.115, 0.219, 0.115, 0.063] 50.0 0.005 33.7 [3.5, 2.0, 3.0, 2, 3, 2.5] 2.667 3.2
i14 Inclusivity 3.111 [0.376, 0.050, 0.341, 0.050, 0.050, 0.133] 33.3 0.003 66.8 [4.0, 3.0, 2.3, 3, 3, 2.8] 3.017 2.5
i14 Aesthetics 2.980 [0.222, 0.211, 0.067, 0.211, 0.222, 0.067] 66.7 0.004 34.6 [4.0, 2.0, 2.7, 2, 4, 2.7] 2.900 2.8
i14 Accessibility 2.607 [0.186, 0.286, 0.148, 0.186, 0.186, 0.009] 16.7 0.002 66.7 [3.0, 2.0, 2.3, 3, 3, 2.6] 2.650 2.4
i13 Practicality 2.550 [0.179, 0.179, 0.219, 0.179, 0.219, 0.025] 83.3 0.002 50.0 [3.0, 3.0, 2.0, 3, 2, 2.5] 2.583 3.0
i13 Inclusivity 1.838 [0.281, 0.072, 0.197, 0.357, 0.072, 0.022] 50.0 0.003 83.4 [2.5, 2.0, 2.3, 1, 2, 1.8] 1.933 2.0
i13 Aesthetics 2.021 [0.014, 0.418, 0.117, 0.400, 0.014, 0.037] 33.3 0.005 66.9 [2.0, 1.0, 2.3, 3, 2, 2.1] 2.067 1.5
i13 Accessibility 2.670 [0.154, 0.154, 0.154, 0.154, 0.315, 0.068] 16.7 0.002 83.7 [3.0, 3.0, 3.0, 3, 2, 2.8] 2.800 2.8
i05 Practicality 2.588 [0.262, 0.173, 0.256, 0.124, 0.124, 0.063] 50.0 0.004 33.8 [3.5, 2.0, 1.7, 3, 3, 2.4] 2.600 1.7
i05 Inclusivity 2.377 [0.217, 0.133, 0.232, 0.217, 0.133, 0.068] 50.0 0.003 66.8 [3.0, 2.0, 1.7, 3, 2, 2.2] 2.317 1.7
i05 Aesthetics 2.716 [0.240, 0.220, 0.181, 0.093, 0.220, 0.045] 66.7 0.004 50.3 [3.5, 2.0, 3.3, 3, 2, 2.6] 2.733 3.3
i05 Accessibility 2.295 [0.100, 0.136, 0.264, 0.315, 0.136, 0.050] 33.3 0.005 83.4 [2.5, 2.0, 1.7, 3, 2, 2.2] 2.233 1.7
i12 Practicality 1.636 [0.192, 0.192, 0.042, 0.192, 0.338, 0.042] 16.7 0.005 83.4 [2.0, 2.0, 1.7, 2, 1, 1.7] 1.733 1.0
i12 Inclusivity 1.591 [0.063, 0.232, 0.068, 0.232, 0.337, 0.068] 16.7 0.006 83.4 [1.5, 2.0, 1.7, 2, 1, 1.7] 1.650 1.0
i12 Aesthetics 1.779 [0.337, 0.107, 0.042, 0.107, 0.365, 0.042] 33.3 0.004 83.4 [2.5, 2.0, 1.7, 2, 1, 1.7] 1.817 1.0
i12 Accessibility 1.974 [0.161, 0.342, 0.112, 0.014, 0.325, 0.046] 50.0 0.004 66.9 [1.5, 3.0, 2.3, 2, 1, 2.1] 1.983 1.0
i20 Practicality 2.893 [0.164, 0.291, 0.237, 0.038, 0.237, 0.034] 66.7 0.006 34.2 [3.5, 4.0, 2.0, 3, 2, 2.8] 2.883 2.6
i20 Inclusivity 1.623 [0.217, 0.133, 0.232, 0.133, 0.217, 0.068] 50.0 0.003 66.8 [1.0, 2.0, 2.3, 2, 1, 1.8] 1.683 2.0
i20 Aesthetics 1.364 [0.192, 0.192, 0.042, 0.338, 0.192, 0.042] 16.7 0.005 83.4 [1.0, 1.0, 1.3, 2, 1, 1.3] 1.267 1.9
i20 Accessibility 2.433 [0.185, 0.302, 0.060, 0.116, 0.278, 0.060] 66.7 0.005 19.1 [1.5, 4.0, 2.7, 3, 1, 2.7] 2.483 2.6
i04 Practicality 2.119 [0.121, 0.350, 0.182, 0.043, 0.274, 0.030] 50.0 0.003 50.2 [2.5, 1.0, 2.7, 2, 3, 2.2] 2.233 3.0
i04 Inclusivity 2.670 [0.154, 0.315, 0.154, 0.154, 0.154, 0.068] 16.7 0.002 83.7 [3.0, 2.0, 3.0, 3, 3, 2.8] 2.800 3.0
i04 Aesthetics 2.169 [0.256, 0.360, 0.044, 0.057, 0.256, 0.027] 50.0 0.004 50.3 [3.0, 1.0, 2.3, 2, 3, 2.1] 2.233 3.0
i04 Accessibility 2.376 [0.264, 0.323, 0.148, 0.148, 0.093, 0.023] 66.7 0.003 33.9 [3.5, 1.0, 3.0, 3, 2, 2.3] 2.467 4.0
i03 Practicality 2.533 [0.313, 0.173, 0.059, 0.173, 0.173, 0.109] 66.7 0.002 16.9 [3.5, 2.0, 2.7, 2, 2, 2.2] 2.400 2.5
i03 Inclusivity 1.733 [0.094, 0.240, 0.313, 0.240, 0.094, 0.021] 50.0 0.006 50.5 [2.0, 1.0, 2.7, 1, 2, 1.7] 1.733 2.8
i03 Aesthetics 1.591 [0.063, 0.232, 0.068, 0.337, 0.232, 0.068] 16.7 0.006 83.4 [1.5, 2.0, 1.7, 1, 2, 1.7] 1.817 1.8
i03 Accessibility 2.550 [0.179, 0.179, 0.179, 0.219, 0.219, 0.025] 33.3 0.002 50.0 [3.0, 3.0, 3.0, 2, 2, 2.5] 2.583 3.0
i15 Practicality 2.779 [0.337, 0.107, 0.042, 0.107, 0.365, 0.042] 33.3 0.004 66.7 [3.5, 2.0, 1.5, 2, 3, 2.7] 2.450 2.5
i15 Inclusivity 1.734 [0.124, 0.124, 0.255, 0.335, 0.124, 0.039] 33.3 0.003 83.5 [2.0, 2.0, 2.3, 1, 2, 1.8] 1.850 2.3
i15 Aesthetics 1.275 [0.149, 0.176, 0.270, 0.176, 0.176, 0.054] 0.0 0.003 66.7 [1.5, 1.0, 1.7, 1, 1, 1.2] 1.233 1.3
i15 Accessibility 2.393 [0.186, 0.286, 0.148, 0.186, 0.186, 0.009] 16.7 0.002 66.7 [2.0, 3.0, 2.7, 2, 2, 2.4] 2.350 2.8